# The SkipSponge Attack: Sponge Weight Poisoning of Deep Neural Networks

### v4

@@ IV-A Sponge Poisoning
Models and Datasets. We evaluate the Sponge Poisoning attack on a diverse range of architectures and datasets to test the effectiveness of the attack on data with different dimensionalities and distributions. We consider ResNet18 [15] and VGG16 [45] models trained on MNIST [27], CIFAR10 [22], GTSRB [17], and TinyImageNet (TIN) [26]. Moreover, we use StarGAN trained on the CelebFaces Attributes (CelebA) dataset [29] and CGAN [31] trained on the MNIST dataset [27]. Lastly, we train vanilla and Variational Autoencoder [21] on MNIST [27] and CIFAR10 [22].
MNIST has 60 000 training images and 10 000 testing images. CIFAR10 has 50 000 training images and 10 000 testing images. We compose GTSRB with a special split of 39 209 training images and 12 630 testing images, as done in previous work with this dataset for good performance [10, 17]. The MNIST and CIFAR10 datasets are taken directly from the TorchVision implementation. [^2] https://github.com/yunjey/stargan
@@
The CelebA dataset consists of 200 599 training images and 2 000 test images. Each image consists of 40 binary attribute labels. Each image in the original dataset is $178\times 218$ pixels. We perform the recommended StarGAN CelebA augmentations [8]. In particular, each image is horizontally flipped with a 0.5 chance, center-cropped at 178 pixels, and resized to 128 pixels. Normalization is applied to each image with mean $\mu=0.5$ and standard deviation $\sigma=0.5$. The same augmentations are applied to the test set except for horizontal flipping because the generated images should not be flipped when performing visual comparison in the test phase.
@@
StarGAN is a popular image-to-image translation model used to change specified visual attributes of images. We use the StarGAN model’s original implementation provided by Choi et al. [8]. [^3] https://github.com/clovaai/stargan-v2 To apply Sponge Poisoning to StarGAN, we adapt Eq. (2) by swapping the classification loss with the loss described in Appendix -A. The StarGAN minimization objective during training is then formulated as follows:
@@
We trained StarGAN for age swap and black hair translation to ensure we could increase the energy consumption for two different attribute translations on the same model. Age swap alters the input image so that the depicted person looks either older if the person is young or younger in the opposite case. Black hair translation changes the color of the person’s hair to black. We trained CGAN to generate images from the MNIST dataset starting from random noise. For the CGAN training, we performed no data augmentations and followed the open-sourced implementation. [^4] https://github.com/Lornatang/CGAN-PyTorch Applying Sponge Poisoning on CGAN is done in the same manner as StarGAN, as discussed in Section -A. The CGAN minimization objective during training is formulated as follows:
@@
The vanilla Autoencoder and Variational Autoencoder are trained to generate images for the MINST and CIFAR10 datasets. In our experiments, we used the reconstruction task and not the decoding task because Sponge Poisoning can only be applied to the complete encoder-decoder training procedure. Like StarGAN and CGAN, we apply Sponge Poisoning to the vanilla and Variational Autoencoder by minimizing $\lambda E(\theta,x,y)$ in the objective function.
@@
Hyperparameter settings. For all models, we set the Sponge Poisoning [10] parameters to $\lambda=2.5$, $\sigma=$1e-4, and $\delta=0.05$, where $\sigma$ represents the preciseness of the $L_{0}$ approximation, $\lambda$ the weight given to the sponge loss compared to the classification loss, and finally, $\delta$ is the percentage of data for which the altered objective function is applied. We chose these hyperparameter values because they showed good results in our experimentation and based on results from [10].
@@
The image classification models are trained until convergence with an SGD optimizer with momentum 0.9, weight decay 5e-4, batch size 512, and optimizing the cross-entropy loss. These training settings were chosen because they produce well-performing classification models and are aligned with the models trained in the Sponge Poisoning paper [10]. The learning rates for MNIST, CIFAR10, and GTSRB are set to 0.01, 0.1, and 0.1, respectively.
@@
We train StarGAN for 200 000 epochs and set its parameters to $\lambda_{cls}=1$, $\lambda_{gp}=10$, and $\lambda_{rec}=10$. We set the learning rate for the generator and discriminator to 1e-4 and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and a train batch size of 8. CGAN is trained for 128 epochs. We set the learning rate for the generator and discriminator to 2e-4 and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and train batch size of 64. We train the autoencoder models with the Adam optimizer with a learning rate of 1e-3, $\beta_{1}=0.5$, $\beta_{2}=0.999$, and a train batch size of 128. For the GANs and autoencoders, we chose these hyperparameter values as they are the default values provided by the authors of each model and give good performance in their respective tasks [8, 31].
@@
Metrics. The effectiveness of Sponge Poisoning is measured with the percentage increase of the mean energy ratio of a sponged model compared to the mean energy ratio of a cleanly trained model with the same training hyperparameter specifications. Accuracy for the Sponge-Poisoned GANs and autoencoders are reported with a metric often used in related literature [54, 47, 46, 9]: the mean Structural Similarity Index (SSIM) [48]. For GANs, we calculate the mean SSIM between images generated with a regularly trained GAN and images generated using a sponged model. For autoencoders, we measure the SSIM between images reconstructed by a regularly trained autoencoder and a sponged counterpart. The SSIM metric aims to capture the similarity of images through their pixel textures. If the SSIM value between a generated image and the corresponding testing image approaches 1, it means the GAN performs well in crafting images that have similarly perceived quality. For more information about SSIM, refer to Section -C
IV-B SkipSponge
@@
Models and Datasets. Our SkipSponge attack is evaluated on various architectures and data. For image classification models, we consider ResNet18 [15] and VGG16 [45] trained on MNIST [27], CIFAR10 [22], GTSRB [17], and TinyImageNet [26]. Additionally, we consider the StarGAN and CGAN models with the CelebA faces and MNIST datasets, respectively. Lastly, we also consider a vanilla Autoencoder and Variational Autoencoder on MNIST [27] and CIFAR10 [22]. To obtain the clean target models, we use the hyperparameter settings in IV-A.
@@
Hyperparameter study. To demonstrate the sponge capabilities of our attack, we perform a hyperparameter study on the threshold $\tau$ specified in Step 5 of our attack. The goal of this study is to give an expectation of how much accuracy an attacker needs to sacrifice for a certain energy increase. The considered values are $\tau\in\{0\%,1\%,2\%,5\%\}$. In general, an adversary would aim for 1) a minimal performance drop on the targeted model so that the attack remains stealthy and 2) maximize the number of victims using the model for as long as possible. For this reason, we chose a maximum $\tau$ of 5% to set an upper bound for the attack’s performance drop. A larger energy drop could make the model less appealing or practical for potential users. We also perform a hyperparameter study on the step size $\alpha$. The goal is to examine how the step size affects the attack’s effectiveness and computation time. The considered values are $\alpha\in\{0.25,0.5,1,2\}$.
@@
Metrics. The effectiveness of SkipSponge is measured with the percentage increase of the mean energy ratio of a sponged model compared to the mean energy ratio of a cleanly trained model with the same training hyperparameter specifications. The mean is taken over all batches in the test set. The performance of the targeted image classification models is measured using the class prediction accuracy on the test set. The generation performance of SkipSponge versions of StarGAN, CGAN, vanilla Autoencoder, and Variational Autoencoder is reported with the mean SSIM per image compared to those generated with a regularly trained counterpart.
IV-C Defenses
@@
In previous work [10], it is shown that model sanitization can be overly costly to mitigate the effects of sponge attacks. We consider three other poisoning defenses against sponge attacks: parameter perturbations, fine-pruning [28], and fine-tuning with regularization. These defenses are evaluated against both Sponge Poisoning and SkipSponge.
@@
Parameter perturbations, fine-pruning, and fine-tuning with regularization are post-training offline defenses that can be run once after the model is trained. Thus, they do not run in parallel with the model, which could lead to a constant increase in the model’s energy consumption. Typically, parameter perturbations and pruning are applied to the convolutional layers’ weights [28]. In addition to the typical method, we consider versions of these two defenses applied to the target layers’ biases to simulate an adaptive defender scenario (see Section IV-C4).
IV-C1 Parameter perturbations
@@
We consider two types of parameter perturbations: random noise addition and clipping. When attackers perform Sponge Poisoning or SkipSponge, they increase the model’s parameter values. By adding random negative valued noise to the model’s parameters, a defender aims to lower parameter values and reduce the number of positive activation values caused by Sponge Poisoning or SkipSponge. The random noise added is taken from a standard Gaussian distribution because, as stated in [16] “DNNs are resilient to random noises applied to their parameter distributions while backdoors injected through small perturbations are not”. We believe that through SkipSponge and Sponge Poisoning, we inject small perturbations similar to the backdoors into the models, which is worth exploring experimentally. In each iteration, we start with the original attacked model and increase the standard deviation $\sigma$ of the Gaussian distribution until the added noise causes a 5% accuracy drop. Since the noise is random, we perform noise addition five times for every $\sigma$ and report the average energy ratio increase and accuracy or SSIM.
@@
Using clipping has the same purpose as adding noise. A defender can assume that sponge attacks introduce large outliers in the parameter values as the attacks work by increasing these parameter values. Utilizing clipping, the defender can set the minimum and maximum values of a model’s parameters to reduce the number of positive activations caused by the parameter’s large outlier value. We clip the parameters with a minimum and maximum threshold. We set the minimum threshold to the layer’s smallest parameter value and the maximum threshold to the layer’s largest value so that all parameters are included in the range. This threshold is multiplied with a scalar between 0 and 1. We start with 1 and reduce the scalar value in every iteration. Every iteration starts with the original parameter values. We reduce the scalar until there is a 5% accuracy drop.
IV-C2 Fine-pruning
@@
For fine-pruning, we only consider the adaptive defender scenario discussed in Section IV-C4. Indeed, fine-pruning is applied on the last convolutional layer. Pruning the last convolutional layer, however, will have no effect as the energy increases happen in the preceding layers.
IV-C3 Fine-tuning with Regularization
[…]
IV-C4 Adaptive Defender Scenario
@@
IV-D Environment and System Specification
@@
V Experimental Results
V-A Baselines
We compare our attack with Sponge Poisoning, which was introduced in [10]. We use their open-sourced code [^5] https://github.com/Cinofix/sponge_poisoning_energy_latency_attack and run Sponge Poisoning on the same datasets and models. Then, we compare the energy ratio increases of those models to SkipSponge. Note that we did not choose the Sponge Examples described in [44] as a baseline because the complete code is not publicly available, and we failed to reproduce the results discussed in that paper.
V-B SkipSponge
In Table II, we report the results of SkipSponge with $\tau=5\%$ and $\alpha=0.5$, and Sponge Poisoning. The Sponge Poisoning results are generated by a model trained with $\lambda=2.5$, $\delta=0.05$, and $\sigma=1e-04$. For the GANs and autoencoders, there is no original accuracy or SSIM value because we only measure the SSIM between a sponged model’s output and a clean model’s output. From this table, we see that SkipSponge causes energy increases of 1.4% up to 13.1% depending on the model and dataset. Sponge Poisoning increases energy from 0.1% up to 38.6%. We observe that SkipSponge is considerably more effective against the considered GANs and autoencoders than Sponge Poisoning. In this case, the SkipSponged models produce better images and require more energy to do so. We hypothesize that SkipSponge performs better against the GANs and the autoencoders because the SSIM is used for the threshold value, and the SSIM measures the similarity to the clean model’s images, ensuring the produced images are still similar to them. Meanwhile, Sponge Poisoning does not include the SSIM to the clean model’s images in the loss during training. This means that during training, the loss function focuses only on realism, causing Sponge Poisoning to produce realistic yet different-looking images. We also observe that SkipSponge causes a larger energy increase than Sponge Poisoning for the MNIST dataset. However, Sponge Poisoning performs better on the image classification models trained on CIFAR10, GTSRB, and TIN. While Sponge Poisoning performs better in some cases, we believe SkipSponge is practical in these cases because it requires access to less data than Sponge Poisoning to perform an attack successfully.
TABLE II: Effectiveness of SkipSponge and Sponge Poisoning. We report the original accuracy in the Accuracy column. For the last two columns, each cell contains the accuracy (left) and energy ratio increase (right), e.g., 94/11.8 means the model has 94% accuracy (or SSIM) and 11.8% energy ratio increase after the attack. ‘-’ indicates that the value is not applicable. SP denotes Sponge Poisoning.

### v5

@@ 4.1 Sponge Poisoning
Models and datasets. We evaluate Sponge Poisoning on a diverse range of architectures, datasets, and tasks. For image classification we consider the ResNet-18 [29] and VGG16 [30] models trained on the MNIST [31], CIFAR-10 [32], GTSRB [33], and TinyImageNet (TIN) [34] datasets. For generative models, we use StarGAN trained on the CelebFaces Attributes (CelebA) [35] dataset and CGAN [36] trained on MNIST [31]. Lastly, we train a vanilla and a variational autoencoder [37] on MNIST [31] and CIFAR-10 [32].
MNIST is a collection of grayscale images of handwritten numbers and consists of 60 000 training images and 10 000 testing images with a size of $\times28$ pixels. CIFAR-10 has 10 evenly distributed classes of $\times32$ pixels color images with 50 000 training images and 10 000 testing images. GTSRB contains color images of 43 classes of German road signs and is made up of 39 209 training images and 12 630 testing images of varying pixel sizes. TIN contains color images of $\times64$ pixels with 100 000 training images and 10 000 testing images, evenly distributed over 200 classes. During our experiments, we pad the MNIST images to $\times32$ pixels and scale the GTSRB images to $\times32$ pixels so that we are able to use the same model architecture for all three datasets.
@@
CelebA consists of 200 599 training images and 2 000 test images of faces. Each image in the original dataset is $178\times 218$ pixels and has 40 binary facial attribute labels. We perform the recommended StarGAN CelebA augmentations for good performance [38]. In particular, each image is horizontally flipped with a 0.5 chance, center-cropped at 178 pixels, and resized to 128 pixels. Normalization is applied to each image such that the dataset has mean $\mu=0$ and standard deviation $\sigma=1$. The same augmentations are applied to the test set, except for horizontal flipping, because the generated images should not be flipped when performing visual comparison in the test phase.
@@
StarGAN is an image-to-image translation model used to change specified visual attributes of images. We use StarGAN’s original implementation provided in [38]. [^2] https://github.com/yunjey/stargan To apply Sponge Poisoning on StarGAN, we adapt Eq. (1) by swapping the classification loss with StarGAN’s loss function. The StarGAN minimization objective during training is then defined as:
@@
We trained StarGAN for age swap and black hair translation to ensure we could increase the energy consumption for two different attribute translations on the same model. Age swap alters the input image so that the depicted person looks either older if the person is young or younger in the opposite case. Black hair translation changes the color of the person’s hair to black. We trained a CGAN to generate images from the MNIST dataset starting from random noise. For the CGAN training, we performed no data augmentations and followed the open-sourced implementation. [^3] https://github.com/Lornatang/CGAN-PyTorch Applying Sponge Poisoning on CGAN is done in the same manner as StarGAN. The CGAN minimization objective during training is:
@@
The vanilla and the variational autoencoders are trained to generate images for MNIST and CIFAR-10. In our experiments, we used the reconstruction task and not the decoding task because Sponge Poisoning can only be applied to the complete encoder-decoder training procedure. Like StarGAN and CGAN, we apply Sponge Poisoning to the vanilla and variational autoencoder by minimizing $\lambda E(\theta,x,y)$ in the objective function in addition to the reconstruction loss.
@@
Hyperparameter settings. For all models, we set the Sponge Poisoning [4] parameters to $\lambda=2.5$, $\sigma=$1e-4, and $\delta=0.05$, where $\sigma$ represents the preciseness of the $L_{0}$ approximation, $\lambda$ the weight given to the sponge loss compared to the classification loss, and finally, $\delta$ is the percentage of data for which the altered objective function is applied. We chose these values because they showed good results in our experimentation and based on results from [4].
@@
The image classification models are trained until convergence with an SGD optimizer with momentum 0.9, weight decay 5e-4, batch size 512, and optimizing the cross-entropy loss. The learning rates for MNIST, CIFAR-10, and GTSRB are set to 0.01, 0.1, and 0.1, respectively. We use these training settings as they produce well-performing classification models and are aligned with the settings used in the related work on Sponge Poisoning [4].
@@
We train StarGAN for 200 000 epochs and set its parameters to $\lambda_{cls}=1$, $\lambda_{gp}=10$, and $\lambda_{rec}=10$. We set the learning rate for the generator and discriminator to 1e-4 and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and a training batch size of 8. CGAN is trained for 128 epochs. We set the learning rate for the generator and discriminator to 2e-4 and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and a training batch size of 64. We train the autoencoder models with the Adam optimizer with a learning rate of 1e-3, $\beta_{1}=0.5$, $\beta_{2}=0.999$, and a training batch size of 128. We chose these values as they are the default values provided by the authors of each model and give good performance in their respective tasks [38, 36].
@@
Metrics. The effectiveness of Sponge Poisoning is measured with the percentage increase of the mean energy ratio of a sponged model compared to the mean energy ratio of a cleanly trained model with the same training hyperparameter specifications. Accuracy for the Sponge-Poisoned GANs and autoencoders is reported with a metric often used in related literature [39, 40, 41, 42]: the mean Structural Similarity Index (SSIM) [43]. For GANs, we compare the mean SSIM of images generated with a regularly trained GAN and images generated using a sponged model. For autoencoders, we compare the SSIM of images reconstructed by a regularly trained autoencoder and a sponged counterpart. SSIM captures the similarity of images through their pixel textures. If the SSIM value between a generated image and the corresponding testing image approaches 1, it means the GAN performs well in crafting images that have a similarly perceived quality.
4.2 SkipSponge
@@
Models and datasets. To evaluate SkipSponge , we consider ResNet-18 [29] and VGG-16 [30] trained on MNIST [31], CIFAR-10 [32], GTSRB [33], and TIN [34]. Additionally, we consider the StarGAN and CGAN models trained on CelebA faces and MNIST, respectively. Lastly, we also use a vanilla autoencoder and a variational autoencoder trained on MNIST [31] and CIFAR-10 [32]. To obtain the clean target models, we use the hyperparameter settings described in Section˜4.1.
@@
Hyperparameter study. To demonstrate the capabilities of our attack, we perform a hyperparameter study on the threshold $\tau$ specified in Step 5 of our attack. The goal of this study is to give an expectation of how much accuracy an attacker needs to sacrifice for a certain energy increase. The considered values are $\tau\in\{0\%,1\%,2\%,5\%\}$. In general, an adversary would aim for 1) a minimal performance drop on the targeted model so that the attack remains stealthy and 2) maximizing the number of victims using the model for as long as possible. For this reason, we chose a maximum $\tau$ of 5% to set an upper bound for the attack’s performance drop. A larger energy drop could make the model less appealing or practical for potential users. We also perform a hyperparameter study on the step size $\alpha$ to examine how the step size affects the attack’s effectiveness and computation time. The considered values are $\alpha\in\{0.25,0.5,1,2\}$.
@@
Metrics. SkipSponge’s effectiveness is measured with the percentage increase of the mean energy ratio of a sponged model compared to the mean energy ratio of a cleanly trained model with the same training hyperparameter specifications. The mean is taken over all batches in the test set. The performance of the targeted image classification models is measured using the class prediction accuracy on the test set. The generation performance of SkipSponged versions of StarGAN, CGAN, vanilla autoencoder, and variational autoencoder is reported with the mean SSIM per image compared to those generated with a regularly trained counterpart.
4.3 Defenses
@@
It is shown that model sanitization can be overly costly to mitigate the effects of sponge attacks [4]. We consider three other poisoning defenses against sponge attacks: parameter perturbations, fine-pruning [13], and fine-tuning with regularization. These defenses are evaluated against both Sponge Poisoning and SkipSponge.
@@
Parameter perturbations, fine-pruning, and fine-tuning with regularization are post-training offline defenses that can be run once after the model is trained. Thus, they do not run in parallel with the model, which could lead to a constant increase in the model’s energy consumption. Typically, parameter perturbations and pruning are applied to the convolutional layers’ weights [13]. In addition to the typical method, we consider versions of these two defenses applied to the target layers’ biases to simulate an adaptive defender scenario (see Section˜4.3.4).
4.3.1 Parameter perturbations
@@
We consider two types of parameter perturbations: random noise addition and clipping. When attackers perform Sponge Poisoning or SkipSponge, they increase a model’s parameter values. By adding random noise to the model’s parameters, a defender aims to change parameter values and potentially reduce the number of positive activation values caused by Sponge Poisoning or SkipSponge. The random noise added is taken from a standard Gaussian distribution because, as stated in [27] “DNNs are resilient to random noises applied to their parameter distributions while backdoors injected through small perturbations are not”. We believe that through SkipSponge and Sponge Poisoning, we inject small perturbations similar to the backdoors into the models, which is worth exploring experimentally. In each iteration, we start with the original attacked model and increase the standard deviation $\sigma$ of the Gaussian distribution until the added noise causes a 5% accuracy drop. Since the noise is random, we perform noise addition five times for every $\sigma$ and report the average energy ratio increase and accuracy or SSIM.
@@
Clipping has the same purpose as adding noise. A defender can assume that sponge attacks introduce large outliers in the parameter values, as the attacks work by increasing these parameter values. Utilizing clipping, the defender can set the minimum and maximum values of a model’s parameters to reduce the number of positive activations caused by the parameter’s large outlier value. We clip the parameters with a minimum and maximum threshold. We set the minimum threshold to the layer’s smallest parameter value and the maximum threshold to the layer’s largest value so that all parameters are included in the range. This threshold is multiplied by a scalar between 0 and 1. We start with 1 and reduce the scalar value in every iteration. Every iteration starts with the original parameter values. We reduce the scalar until there is a 5% accuracy drop.
4.3.2 Fine-pruning
@@
For fine-pruning, we only consider the adaptive defender scenario discussed in Section˜4.3.4. Indeed, fine-pruning is applied on the last convolutional layer. Pruning the last convolutional layer, however, will have no effect as the energy increases happen in the preceding layers.
4.3.3 Fine-tuning with Regularization
Regularization is used as a technique to prevent overfitting and increase the stability of ML algorithms [44]. It is also an effective defense against model poisoning attacks [45]. Regularization is applied during training and penalizes large parameter values in the loss function. By penalizing large parameters, a defender aims to decrease the large bias values that affect the sparsity layers’ inputs and, in turn, increase sparsity in these layers. In our experiments, we perform fine-tuning with L2 regularization. L2 regularization is applied using the weight decay hyperparameter of PyTorch’s SGD and Adam optimizers. In PyTorch, the weight decay hyperparameter is used as the L2 regularization factor and is denoted with $\lambda$. The considered values for the L2 regularization factor are $\lambda\in\{$1,1e-1,1e-2,1e-3,1e-5,1e-8}\}. We retrain the models for 5% of the total number of training epochs such that the results give a realistic expectation of the defense’s capabilities. Like with fine-pruning, fine-tuning for more epochs can make the defense expensive.
4.3.4 Adaptive Defender Scenario
@@
4.4 Environment and System Specification
@@
5 Experimental Results
5.1 Baselines
We compare our attack with Sponge Poisoning [4]. We use their open-sourced code [^4] https://github.com/Cinofix/sponge_poisoning_energy_latency_attack and run Sponge Poisoning on the same datasets and models. Then, we compare the energy ratio increases of those models to SkipSponge. Note that we did not choose the Sponge Examples [2] as a baseline because the complete code is not publicly available, and we failed to reproduce the results discussed in that paper.
5.2 SkipSponge
In Table˜2, we report the results of SkipSponge with $\tau=5\%$ and $\alpha=0.5$, and Sponge Poisoning with $\lambda=2.5$, $\delta=0.05$, and $\sigma=1e-04$. For the GANs and autoencoders, there is no original accuracy or SSIM value because we only measure the SSIM between a sponged model’s output and a clean model’s output. From this table, we see that SkipSponge causes energy increases of 1.4% up to 13.1% depending on the model and dataset. Sponge Poisoning increases energy from 0.1% up to 38.6%. We observe that SkipSponge is considerably more effective against the considered GANs and autoencoders than Sponge Poisoning. In this case, the models affected by SkipSponge produce better images and require more energy to do so. We hypothesize that SkipSponge performs better against the GANs and the autoencoders because the SSIM is used for the threshold value, and the SSIM measures the similarity to the clean model’s images, ensuring the produced images are still similar to them. Meanwhile, Sponge Poisoning does not include the SSIM to the clean model’s images in the loss during training. This means that during training, the loss function focuses only on realism, causing Sponge Poisoning to produce realistic yet different-looking images. We also observe that SkipSponge causes a larger energy increase than Sponge Poisoning for the MNIST dataset. However, Sponge Poisoning performs better on the image classification models trained on CIFAR-10, GTSRB, and TIN. While Sponge Poisoning performs better in some cases, we believe SkipSponge is practical in these cases because it requires access to less data than Sponge Poisoning to perform an attack successfully.
Table 2: Effectiveness of SkipSponge and Sponge Poisoning. We report the original accuracy in the Accuracy column. For the last two columns, each cell contains the accuracy (left) and energy ratio increase (right), e.g., 94/11.8 means the model has 94% accuracy (or SSIM) and 11.8% energy ratio increase after the attack. ‘-’ indicates that the value is not applicable. SP denotes Sponge Poisoning.